OpenStudy (anonymous):

An orchestra consists of 25 members. The youngest member is 32 years old. She leaves and is replaced by a new member who is 30 years old.By how much does the replacement change the average age of the members of the orchestra?

OpenStudy (anikhalder):

hmm...lets's see it like this:- You have 5 apples...4 of them cost 5$ each whereas the 5th apple costs 10$.. So total cost of the 5 apples=(4*5)+10=30$ So, average cost(let's take this as initial average cost)=30/5=6$ Now,I take away your 5th apple(the 10$ apple) and replace it with a 5$ apple.. So, now you have 5 apples each with a price of 5$ Total cost f the 5 apples you have now=(5*5)=25$ Hence the average cost now(new average cost)=25/5=5$ So the new average cost differs from the initial average cost by=6-5=1$ New average cost is 1$ less than the initial average cost.. Hope you can use this concept to nail your problem..:)